With Washimi, Hirata, Aso and Satou
Bottom Filter readout
We realized the strange 30 mHz oscillation observed in the BF and F1 is aliasing noise downconverted when seeing the slow channel. In the fast channel it's high frequency noise.
- A low pass filter removes the high frequency noise
- The unusual feature is that the amplitude of the noise changes with the LVDT position. This points out to the LVDT card as the origin of the noise, but we need to check more.
Mechanics of the Payload and the Bottom Filter
- In the morning we locked the BF sides to the security structure and pulled the BF-IM suspension rod with the hand just to check how far down it reached. It went down only to about 35 um, which is too high. It was definitely hiting a hard stop.
- Given that the stop looed like a hard one, we proceeded to withdraw higher the lock nuts of the keystones, in order to give more space in case any of them was creating the problem. We used a nut driver size 10 for a M6 nut. Fortunately it worked: after the adjustment we were able to move the keystone down to -1.2 mm.
- I still don't know what determines this distance, but in SR3 it is -1.470 mm, which is similar.
- In the afternoon Hirata-san withdrew the nuts even more but the keystone didn't move further down.
- I checked the 3D-CAD and, by design we should be able to move it down by -1.75 mm, the limiting component is the OSEM flag whose base attaches around the magnet. This means -1.470 mm and -1.2 mm are alright.
- Our aim was to instal the limit switches for the Fishing Rod (FR), but we realized that the holes we need to put M4 screws into are not threaded at the top side of the IP table, but onlhy at the bottom. We came up with an alternative plan which requires inserting long M3 screws from the bottom and using an M3 washer with an O.D less than 8 mm. The importance of using a washer of such a size is that we're using those holes to hold a cable clamp base. The details will be incorporated into the 3D CAD soon.
- Despite the tempoarary mechanical drawback we installed the limit switch trigger piece, which mounts on a component that is already there.
- We also came up with a plan for connecting the limit switches cables to the D-sub 9 connector underneath the IP table. Tomorrow we will proceed connecting the cables and testing them using the medm screen.